OpenStudy (anonymous):

simplify by removing factors of 1..... (w^2-81)/(9-w)

OpenStudy (saifoo.khan):

Factorize.

OpenStudy (saifoo.khan):

\[\frac{(w-9)(w+9)}{9-w}\]Can you solve it now?

OpenStudy (anonymous):

I tired just not sure its right -w-9?

OpenStudy (saifoo.khan):

\[\frac{(w-9)(w+9)}{-1(-9+w)}\]

OpenStudy (saifoo.khan):

\[\frac{\cancel{(w-9)}(w+9)}{-1\cancel{(-9+w)}}\]

OpenStudy (saifoo.khan):

\[\frac{w+9}{-1}=-w-9\]

OpenStudy (anonymous):

ty :)

OpenStudy (saifoo.khan):

uw! :0

OpenStudy (saifoo.khan):

:)